QUESTION 4 a) only have RMIL.167 now. At what rate must your RMI1,167 be compounded anmualy for In 10 years, you had really like to have RM20,000 to buy a new Perodus Asia, but you now. At what rate must your RM11,167 be compounded annually for it to grow to RM20,000 in 10 years? b) Sup you are investing RM5.000 at the end ofeach year in account that pays 7%. How c) Fnd the effective annual rate (EAR) for a 8% nominal annual rate when interest is d) Whats the present value of the following uneven cash flow stream: RM400 in Year 1 e) Nozita takes out a loan for RM30,000 at 10% interest per year and is required to pay of long will it be before your account is worth RM51,300 compounded quarterly? RMS00 in Year 2, RM500 in Year 3, and RM500 in Year 4 ifthe interest rate is 10%? this loan with five equal annual payments. What is the required payment? loan of RM6,000 at 15% is to be amortized in three years. Required: i. Calculate dt yearly paymént and shows the ii. Construct an amortization schedule and fill up the table below 1 A of interest in the first year (4 marks) (8 marks) Amount owed on the Annuity Repayment Ending principal at payment (RM) principal (RM) of Year | the beginning | principalbalance ·(RM) of year (RM) (RM PV uneven CLCF2CFn
Answer #1
a) We have the equality
20000 = 11167*(1+r)^10, where r is the required rate.
Solving for r
r = (20000/11167)^(1/10)-1 = 6.00%
b) We have the equality
51300 = 5000*PVIFA(7,n), where n = the required number of
years.
PVIFA(7,n) = 51300/5000 = 10.2600
From the interest factor tables
interest factor at 7% for n = 18 =10.0591 and for n = 19 it is
10.3356
By simple interpolation, the value of n for factor of 10.2600
= 18+(10.26-10.0591)/(10.3356-10.0591) = 18.73
c) EAR = (1+r/m)^m-1, where m is the number of compounding
done per year.
EAR = 1.02^4-1 = 8.24%
d) PV = 400/1.1+800/1.1^2+500/1.1^3+500/1.1^4 = $    1,741.96
e) Required annual payment (using the formula for loan
amortization) = 30000*0.10*1.1^5/(1.1^5-1) = $    7,913.92
f) i) Required annual payment (using the formula for loan
amortization) = 6000*0.15*1.15^3/(1.15^3-1) = $    2,627.86
Interest in the first year = 6000*15% = $       900.00
ii)
Year Amount owed on the principal at the beginning of the year Annuity Payment Interest Repayment of Principal Ending balance
1 $                                                                                               6,000.00 $    2,627.86 $             900.00 $ 1,727.86 $   4,272.14
2 $                                                                                               4,272.14 $    2,627.86 $             640.82 $ 1,987.04 $   2,285.10
3 $                                                                                               2,285.10 $    2,627.86 $             342.76 $ 2,285.10 $         -0.00